# Scalable multiphoton generation from cavity-synchronized single-photon sources

### v2

@@ III.2 Efficiency of single-photon generation
A nearly deterministic single-photon source should have ${\cal P}_{1}\approx 1$, other few-photon probabilities ${\cal P}_{0}$ and ${\cal P}_{2}$ should be strongly suppressed, and ${\cal P}_{n\geq 3}\approx 0$ negligible. This also applies to each of the probabilities ${\cal P}_{n}^{j}$ in our synchronized setup, if we want it to operate as a collection of $N$ independent emitters.
To verify this, we have computed the few-photon probabilities ${\cal P}_{0}$ (green), $1-{\cal P}_{1}$ (blue), and ${\cal P}_{2}$ (brown) for a setup with $N=2$ superconducting qubits, varying different parameters around the set C from Table 1. The results in Figs. 3(a)-(f), have been computed using the full model (6) and the effective equation (12), and are shown in lines and markers, respectively. The agreement between both methods confirms that the adiabatic elimination of the cavity is an excellent approximation over a broad range of parameters.
@@
In Fig. 3(a) we plot the effect of cavity occupation $|\alpha_{\rm ss}|^{2}$ (controlled via the cavity drive $\Omega_{0}$) on the single-photon efficiency of the SPS. All errors reduce with the cavity population $|\alpha_{\rm ss}|^{2},$ because a larger cavity-mediated driving $\Omega_{\rm cm}=|\alpha_{\rm ss}|g$ increases the speed at which the qubit is excited and reduces the probability of emitting two consecutive photons. However, the driving strength is limited by our need to have an off-resonant cavity and by the anharmonicity of the qubits [cf. Sec. III.1].
@@
Fig. 3(b) illustrates the need to find an optimum value of the coupling $g$, within inequalities in Eq. (18). Initially, increasing $g$ increases the driving $\Omega_{\rm cm}$ and reduces errors. However, a large coupling strength enhances the cavity-mediated dipole-dipole interactions $J_{\rm cm}\sim g^{2}$ and the collective decay $\gamma_{\rm cm}\sim g^{2}$. Both result in the transfer of excitations between qubits, increasing the probability of no emission ${\cal P}_{0}$ and of two photon emission ${\cal P}_{2}.$ For the parameters in Table 1 we find an optimal operation point around $g/\kappa=4\cdot 10^{-4}$.
@@
In Fig. 3(c) we analyze the effect of the coupling to the antenna coupling $\gamma.$ Once more, there is an optimal point that satisfies (18), balancing the imperfections due to unwanted re-excitations ($\Omega_{\rm cm}\gg\gamma$) and cavity-mediated effects ($\gamma\ll J_{\rm cm},\gamma_{\rm cm}$). The optimal lays around $\gamma/\Omega_{\rm cm}\sim 2\cdot 10^{-2}$ for the parameter set C in Table 1.
@@
Fig. 3(d) clearly shows an optimal operation of the emitters when they are on resonance with the drive $\delta=\delta_{\rm cm}$. At this point, a $\pi$ pulse excites the qubits with high fidelity, and the Lamb-shift is compensated.
@@
Finally, in Fig. 3(e) and Fig. 3(f), we observe that the dephasing $\gamma_{\phi}$ and the losses into uncontrolled channels $\gamma_{\rm loss}$ have a negligible influence on the few-photon statistics ${\cal P}_{n}$ as long as they satisfy $\gamma_{\phi},\gamma_{\rm loss}\ll\gamma$ [cf. Eq. (18)]. Nevertheless, when $\gamma_{\phi},\gamma_{\rm loss}\sim\gamma$ we see a rapid increase of events where no photon is detected ${\cal P}_{0}.$ This is due to emitters becoming effectively off-resonant from the drive when $\gamma_{\phi}\sim\gamma$, or due to a decrease in the the collection efficiency in the antenna, when $\gamma_{\rm loss}\sim\gamma.$
Figure 4: Multi-photon contamination of the SPS measured via the second-order correlation function. (a) Schematic of the Hanbury Brown and Twiss (HBT) experiment using a beam-splitter (BS) and two-photon coincidence measurements. (b) HBT correlation function $G^{(2)}_{\rm HBT}(\tau)$ as a function of delay $\tau$, in the case of a train of 4 excitation pulses with repetition rate $R=\gamma/15$. Parameters correspond to set C of Table 1. (c) Enlarged plot of $G^{(2)}_{\rm HBT}(\tau)$ at the zeroth peak, $\tau\sim 0$, for three values of couplings $g/\kappa=[1,5,10]\cdot 10^{-4}$. (d,e) Normalized correlation function $g^{(2)}_{\rm HBT}[0]$ at zero delay, as function of dephasing $\gamma_{\phi}/\gamma$ and out-coupling inefficiency $\gamma_{\rm loss}/\gamma$, respectively. Other parameters correspond to set C (blue) and set B (red). Decoherence has a small detrimental effect on the observed multi-photon contamitation via $g^{(2)}_{\rm HBT}[0]$.
@@ III.3 Efficiency of synchronized N-photon generation
To quantify more precisely the independence of the $N$ emitted photons, we define the $N$-photon dependence or demultiplexing error as,
@@
which describes the deviation from the ideal limit of $N$ perfectly independent and synchronized SPSs. This quantity is also related to the demultiplexing inefficiency in Eq. (2) as $S_{N}=1+D_{N}$.
We have computed $P_{N}$ and $D_{N}$ in the simplest case of two SPSs, using the same photon counting methods introduced in Appendix B. The results of these simulations are shown as red curves of Figs. 3(a)-(f) [cf. red curves and right vertical axis]. We observe that $D_{2}$ is nearly insensitive to changes in the cavity occupation $|\alpha_{\rm ss}|^{2}$, the dephasing $\gamma_{\phi}/\gamma$, and out-coupling efficiency $\gamma_{\rm loss}/\gamma$, with typical values on order $D_{2}\sim 10^{-7}$ for parameters around set C of Table 1. Nevertheless, the variables $g$, $\gamma$, and $\delta$ can drastically change $D_{2}$ in various orders of magnitude. In particular, for larger cavity coupling $g$ or smaller antenna decay $\gamma$, $D_{2}$ increases because the cavity-mediated correlations $\sim g^{2}$ become more important in the timescale of the photon emission $\sim 15/\gamma$ [cf. right inequality in Eq. (18)].
Figure 5: Indistinguishability of single-photons from two different SPSs quantified by Hong-Ou-Mandel (HOM) interference. (a) Schematic of the HOM setup using a beam splitter to interfere two nearly independent photons, followed by two-photon coincidence measurements. (b) HOM correlation function $G^{(2)}_{\rm HOM}(\tau)$ as a function of delay $\tau$, in the case of a train of 4 excitation pulses with repetition rate $R=\gamma/15$. Parameters correspond to set C of Table 1. (c) Enlarged plot of $G^{(2)}_{\rm HOM}(\tau)$ at the zeroth peak, $\tau\sim 0$, for three values of couplings $g/\kappa=[1,5,10]\cdot 10^{-4}$. (d,e) Normalized correlation function $g^{(2)}_{\rm HOM}[0]$ at zero delay, as function of dephasing $\gamma_{\phi}/\gamma$ and out-coupling inefficiency $\gamma_{\rm loss}/\gamma$, respectively. Other parameters correspond to set C (blue) and set B (red). The indistinguishability dramatically reduces with dephasing, a behaviour in good agreement with the analytical prediction $g^{(2)}_{\rm HOM}[0]\approx(\gamma_{\phi}/\gamma)/(1+\gamma_{\phi}/\gamma)$ Bylander et al. (2003).
@@ III.4 Single-photon purity
where $\tau$ is the time delay between the two photon detections, and $b^{j}_{\rm out}(t)$ is the annihilation operator for an outgoing photon on the $j$-th antenna Gardiner and Zoller (2004). Appendix C contains details on the input-output theory and the calculation of these correlations. For simplicity, our analysis assumes a homogeneous setup and thus we omit the $j$ index in $G_{\rm HBT}^{(2)}(\tau)$.
In Fig. 4(b) we show the behaviour of the correlation $G_{\rm HBT}^{(2)}(\tau)$ in our setup for train of 4 excitation pulses with a repetition rate
@@
We observe clear peaks at integer multiples of the repetition time $\tau=n/R$, corresponding to the detection of two photons coming from different pulses. Relevant to characterize the few-photon statistics of the SPS is the small peak that appears near zero time delay, $\tau\sim 0.$ A small area in this peak signals a small probability of two- and multi-photon emission Kiraz et al. (2004). In Fig. 4(c) we enlarge the region $\tau\sim 0$ and show that the area of the zeroth peak increases with $g/\kappa,$ consistently with the behavior of ${\cal P}_{2}$ in Fig. 3(b) for the same parameters.
@@
Here, the numerator corresponds to the area of the zeroth peak at $\tau\sim 0$ and the normalization is the area of one of the high peaks at $\tau\sim n/R$, and thus $g_{\rm HBT}^{(2)}[0]\sim{\cal P}_{2}/({\cal P}_{1})^{2}$. In Figs. 4(d)-(e), we plot the normalized correlation $g_{\rm HBT}^{(2)}[0]$ as a function of $\gamma_{\phi}/\gamma$ and $\gamma_{\rm loss}/\gamma$, respectively. We observe that $g_{\rm HBT}^{(2)}[0]$ is very insensitive to both types of decoherence, which is also consistent to the behaviour of ${\cal P}_{2}$ in Fig. 3(e)-(f) for the same parameters. In addition, we see that parameter set B gives a two-photon contamination of $g_{\rm HBT}^{(2)}[0]\sim 2.5\cdot 10^{-3}$, slightly lower than parameter set C [cf. Fig. 4(e)-(f) and Table 1]. Finally, we note that the single-photon purity $p$ is sometimes defined as $p=1-g_{\rm HBT}^{(2)}[0]$, which for parameter sets B and C of our circuit-QED implementation is higher than $p=0.99$.
@@ III.5 Indistinguishability
Another important aspect of a high-performance multiphoton demultiplexing scheme is the degree of indistiguishability of the photons emitted by two different sources Senellart et al. (2017). Experimentally, the distinguishability of two sources is typically quantified using Hong-Ou-Mandel (HOM) interference as shown in Fig. 5(a). Photons coming from two sources interfere on a beam-splitter (BS), and a pair of detectors count coincidences. The result of these measurements are the HOM correlations $G_{\rm HOM}^{(2)}(\tau)$, which for pulsed emission read Fischer et al. (2016)
@@
where $\tau$ is the delay time between the two photon detections. In addition, $b_{BS}^{1}(t)=[b^{1}_{\rm out}(t)+b^{2}_{\rm out}(t)]/\sqrt{2}$, and $b_{BS}^{2}(t)=[b^{1}_{\rm out}(t)-b^{2}_{\rm out}(t)]/\sqrt{2}$, describe the output fields of photons after interfering at the BS [cf. Appendix C].
@@
Figure 6: Scalability performance of the synchronized multiphoton generation scheme. (a) NN-photon generation efficiency $P_{N}$ as a function of $N$, for parameter sets A (green), B (blue), C (red), and D (grey), as given in Table 1. Data points up to $N=12$ correspond to QT calculations averaged over $M=6000$ trajectories, which show excellent agreement (within the statistical error $\Delta P_{N}\sim 10^{-3}$) with the prediction for perfectly independent and synchronized SPSs, i.e. $P_{N}\approx({\cal P}_{1})^{N}$. [cf. black lines and inset]. (b) NN-photon generation rate $C_{N}=RP_{N}$ for pulsed excitation with repetition rate $R=\gamma/15$. Despite parameter set D shows the lowest $P_{N}$ efficiency, its faster repetition rate $R\sim 1.2$MHz allows for appreciably higher $C_{N}$. (c) Demultiplexing error $D_{N}$ up to $N=5$, and quadratic fit $D_{N}=\epsilon N(N-1)$ (cf. solid lines and inset). The fitted scaling factor depends on the system parameters [cf. Eq. (28)] and read $\epsilon_{A}=6.7\times 10^{-7}$, $\epsilon_{B}=6.7\times 10^{-7},\epsilon_{C}=1.0\times 10^{-7}$, and $\epsilon_{D}=1.3\times 10^{-8}$ for parameter sets A, B, C, and D, respectively. The extrapolation of all these calculations are valid as long as $D_{N}\lesssim 1$.
@@
In analogy to Eq. (24), the numerator of Eq. (26) corresponds to the area below the zeroth peak and the normalization to the area of one of the high peaks at the repetition times $\tau=n/R$. The indistinguishability of two SPSs can be simply defined as $I=1-g^{(2)}_{\rm HOM}[0].$
We have computed $G_{\rm HOM}^{(2)}(\tau)$ as a function of $\tau$, for a setup with two emitters and a train of 4 excitation pulses [cf. Fig. 5(b)]. Here, we observe clear peaks at the repetition times $\tau=n/R$ due to the detection of two nearly independent photons coming from two different pulses. The correlations $G_{\rm HOM}^{(2)}(\tau)$ are strongly suppressed at the origin, $\tau\sim 0$, showing only a minor zeroth peak which is enlarged in Fig. 5(c). The small area of this zeroth peak manifests the high indistinguishability of the single-photons prepared in our setup with parameter set C. We also see that this behaviour is nearly insensitive to the coupling $g/\kappa$, but may strongly depend on the decoherence parameters.
Figures 5(d)-(e) display $g^{(2)}_{\rm HOM}[0]$ as function of the decoherence rates $\gamma_{\phi}/\gamma$ and $\gamma_{\rm loss}/\gamma.$ Note that in our implementation with parameters B and C we reach values above $I=0.99$, as inferred from Fig. 5(e) and Table 1. We also find that the distinguishability $g^{(2)}_{\rm HOM}[0]$ is quite insensitive to $\gamma_{\rm loss}/\gamma$, but it dramatically increases with dephasing, following the analytical prediction $g^{(2)}_{\rm HOM}[0]\approx(\gamma_{\phi}/\gamma)/(1+\gamma_{\phi}/\gamma)$ Bylander et al. (2003). This contrasts with the behaviour of $D_{2}$ in Fig. 3(e), illustrating that the dependence error $D_{2}$ of two SPSs is qualitatively different from the distinguishability $g^{(2)}_{\rm HOM}[0]$ measured by HOM interference. Consequently, both figures of merit have to be considered when designing high-performance multiplexed and synchronized SPSs.

### v3

@@ III.2 Efficiency of single-photon generation
A nearly deterministic single-photon source should have ${\cal P}_{1}\approx 1$, other few-photon probabilities ${\cal P}_{0}$ and ${\cal P}_{2}$ should be strongly suppressed, and ${\cal P}_{n\geq 3}\approx 0$ negligible. This also applies to each of the probabilities ${\cal P}_{n}^{j}$ in our synchronized setup, if we want it to operate as a collection of $N$ independent emitters. To verify this, in Figs. 3(a)-(f) we have computed the few-photon probabilities ${\cal P}_{0}$ (green), $1-{\cal P}_{1}$ (blue), and ${\cal P}_{2}$ (brown) for a setup with $N=2$ superconducting qubits, varying different parameters around the set A of Table 1.
@@
In Fig. 3(a) we plot the effect of cavity occupation $|\alpha_{\rm ss}|^{2}$ (controlled via the cavity drive $\Omega_{0}$) on the single-photon efficiency of the SPS. All errors reduce or at least stay nearly constant when increasing the cavity population $|\alpha_{\rm ss}|^{2}$, because a larger cavity-mediated driving $\Omega_{\rm cm}=|\alpha_{\rm ss}|g$ increases the speed at which the qubit is excited and reduces the probability of emitting two consecutive photons. However, the driving strength is limited by our need to have an off-resonant cavity and by the anharmonicity of the qubits [cf. Sec. III.1].
@@
Fig. 3(b) illustrates the need to find an optimum value of the coupling $g$, within inequalities in Eq. (19). Initially, increasing $g$ increases the driving $\Omega_{\rm cm}$ and reduces errors. However, a large coupling strength enhances the cavity-mediated dipole-dipole interactions $J_{\rm cm}\sim g^{2}$ and the collective decay $\gamma_{\rm cm}\sim g^{2}$. Both result in the transfer of excitations between qubits, increasing the probability of no emission ${\cal P}_{0}$ and two-photon emission ${\cal P}_{2}.$ For the parameters A in Table 1 we find an optimal operation point around $g/\kappa={\color[rgb]{0,0,0}10^{-1}}$.
@@
In Fig. 3(c) we analyze the effect of the coupling to the antenna coupling $\gamma.$ Once more, there is an optimal point that satisfies (19), balancing the imperfections due to unwanted re-excitations ($\Omega_{\rm cm}\gg\gamma$) and cavity-mediated effects ($\gamma\ll J_{\rm cm},\gamma_{\rm cm}$). The minima lay around $\gamma/\Omega_{\rm cm}\sim[10^{-2},10^{-1}]$ for the parameter set A in Table 1.
@@
Fig. 3(d) shows an optimal operation of the emitters when they are on resonance with the drive $\delta\sim\delta_{\rm cm}$. At this point, a $\pi$ pulse excites the qubits with high fidelity, and the Lamb shift is compensated, which is the main qubit’s frequency shift in our setup.
@@
Finally, in Fig. 3(e) and Fig. 3(f), we observe that the dephasing $\gamma_{\phi}$ and the losses into uncontrolled channels $\gamma_{\rm loss}$ have a negligible influence on the few-photon statistics ${\cal P}_{n}$ as long as they satisfy $\gamma_{\phi},\gamma_{\rm loss}\ll\gamma$ [cf. Eq. (19)]. Nevertheless, when $\gamma_{\phi},\gamma_{\rm loss}\sim\gamma$ we see a rapid increase of events where no photon is detected ${\cal P}_{0}.$ This is due to emitters becoming effectively off-resonant from the drive when $\gamma_{\phi}\sim\gamma$, or due to a decrease in the collection efficiency in the antenna, when $\gamma_{\rm loss}\sim\gamma.$
@@ III.3 Efficiency of synchronized N-photon generation
To quantify more precisely the independence of the $N$ emitted photons, we compute the $N$-photon dependence or demultiplexing error $D_{N}$ from the definition in Eq. (3) as,
@@
which describes the deviation from the ideal limit of $N$ perfectly independent and synchronized SPSs.
We have computed $P_{N}$ and $D_{N}$ in the simplest case of two SPSs as an example, using the same photon counting methods introduced in Appendix B. The results of these simulations are shown as red curves of Figs. 3(a)-(f) [cf. red curves and right vertical axis]. We observe that $D_{2}$ is nearly insensitive to changes in the cavity occupation $|\alpha_{\rm ss}|^{2}$, the detuning $\delta$, the dephasing $\gamma_{\phi}/\gamma$, and out-coupling efficiency $\gamma_{\rm loss}/\gamma$, with typical values on order $D_{2}\sim{\color[rgb]{0,0,0}10^{-4}}$ for parameters around set A of Table 1. Nevertheless, the variables $g$ and $\gamma$ can drastically change $D_{2}$ in various orders of magnitude. In particular, for larger cavity coupling $g$ or smaller antenna decay $\gamma$, $D_{2}$ increases because the cavity-mediated correlations $\sim g^{2}$ become more important in the timescale of the photon emission $\sim 15/\gamma$ [cf. right inequality in Eq. (19)].
@@
Figure 4: Multi-photon contamination of the SPS measured via the second-order correlation function. (a) Schematic of the Hanbury Brown and Twiss (HBT) experiment using a beam-splitter (BS) and two-photon coincidence measurements. (b) HBT correlation function $G^{(2)}_{\rm HBT}(\tau)$ as a function of delay $\tau$, in the case of a train of 2 excitation pulses with repetition rate $R=\gamma/15$. Parameters correspond to set A of Table 1. (c) Enlarged plot of $G^{(2)}_{\rm HBT}(\tau)$ at the zeroth peak, $\tau\sim 0$, for three values of couplings $g/\kappa={\color[rgb]{0,0,0}[2,8,20]\cdot 10^{-2}}$. (d,e) Normalized correlation function $g^{(2)}_{\rm HBT}[0]$ at zero delay, as function of dephasing $\gamma_{\phi}/\gamma$ and out-coupling inefficiency $\gamma_{\rm loss}/\gamma$, respectively. Decoherence has a small detrimental effect on the observed multi-photon contamination via $g^{(2)}_{\rm HBT}[0]$.
@@ III.4 Single-photon purity
where $\tau$ is the time delay between the two-photon detections, and $b^{j}_{\rm out}(t)$ is the annihilation operator for an outgoing photon on the $j$-th antenna Gardiner and Zoller (2004). Appendix C contains details on the input-output theory and the calculation of these correlations. For simplicity, our analysis assumes a homogeneous setup and thus we omit the $j$ index in $G_{\rm HBT}^{(2)}(\tau)$.
In Fig. 4(b) we show the behaviour of the correlation $G_{\rm HBT}^{(2)}(\tau)$ in our setup for parameters A and a train of 2 consecutive excitation pulses with a repetition rate
@@
We observe a clear peak at the repetition time $\tau=1/R$, corresponding to the detection of two photons coming from different pulses. Relevant to characterize the few-photon statistics of the SPS is the small peak that appears near zero time delay, $\tau\sim 0.$ A small area in this peak signals a small probability of two- and multi-photon emission Kiraz et al. (2004). In Fig. 4(c) we enlarge the region $\tau\sim 0$ and show that the area of the zeroth peak increases with $g/\kappa,$ consistently with the behavior of ${\cal P}_{2}$ in Fig. 3(b) for the same parameters.
@@
Here, the numerator corresponds to the area of the zeroth peak at $\tau\sim 0$ and the normalization is the area of the high peak at $\tau=1/R$, and thus $g_{\rm HBT}^{(2)}[0]\sim{\cal P}_{2}/({\cal P}_{1})^{2}$. In Figs. 4(d)-(e), we plot the normalized correlation $g_{\rm HBT}^{(2)}[0]$ as a function of $\gamma_{\phi}/\gamma$ and $\gamma_{\rm loss}/\gamma$, respectively. We observe that $g_{\rm HBT}^{(2)}[0]$ is very insensitive to both types of decoherence, which is also consistent with the behavior of ${\cal P}_{2}$ in Fig. 3(e)-(f) for the same parameters. For parameters A, we obtain a two-photon contamination of $g_{\rm HBT}^{(2)}[0]{\color[rgb]{0,0,0}\sim 1.4\cdot 10^{-2}}$. We also calculated this quantity for parameters B-D of Table 1, for which we obtained $g_{\rm HBT}^{(2)}[0]\sim 2.8\cdot 10^{-2}$, $g_{\rm HBT}^{(2)}[0]\sim 4.4\cdot 10^{-3}$, and $g_{\rm HBT}^{(2)}[0]\sim 1.3\cdot 10^{-2}$, respectively. Finally, we note that the single-photon purity $p$ is typically defined as $p=1-g_{\rm HBT}^{(2)}[0]$, which for parameter sets A, B, C, and D read $p=0.986$, $p=0.972$, $p=0.996$, and $p=0.987$, respectively. We see that slower sources have a larger single-photon purity (see A and C compared to B and D).
@@ III.5 Indistinguishability
Figure 5: Indistinguishability of single-photons from two different SPSs quantified by Hong-Ou-Mandel (HOM) interference. (a) Schematic of the HOM setup using a beam splitter to interfere two nearly independent photons, followed by two-photon coincidence measurements. (b) HOM correlation function $G^{(2)}_{\rm HOM}(\tau)$ as a function of delay $\tau$, in the case of a train of 2 excitation pulses with repetition rate $R=\gamma/15$. Parameters correspond to set A of Table 1. (c) Enlarged plot of $G^{(2)}_{\rm HOM}(\tau)$ at the zeroth peak, $\tau\sim 0$, for three values of couplings $g/\kappa={\color[rgb]{0,0,0}[2,8,20]\cdot 10^{-2}}$. (d,e) Normalized correlation function $g^{(2)}_{\rm HOM}[0]$ at zero delay, as function of dephasing $\gamma_{\phi}/\gamma$ and out-coupling inefficiency $\gamma_{\rm loss}/\gamma$, respectively. The indistinguishability dramatically reduces with dephasing, a behavior in good agreement with the analytical prediction $g^{(2)}_{\rm HOM}[0]\approx({\color[rgb]{0,0,0}2}\gamma_{\phi}/\gamma)/(1+{\color[rgb]{0,0,0}2}\gamma_{\phi}/\gamma)$ Bylander et al. (2003).
Another important aspect of a high-performance multiphoton demultiplexing scheme is the degree of indistinguishability of the photons emitted by two different sources Senellart et al. (2017). Experimentally, the distinguishability of two sources is typically quantified using Hong-Ou-Mandel (HOM) interference as shown in Fig. 5(a). Photons coming from two sources interfere on a beam-splitter (BS), and a pair of detectors count coincidences. The result of these measurements are the HOM correlations $G_{\rm HOM}^{(2)}(\tau)$, which for pulsed emission read Fischer et al. (2016)
@@
where $\tau$ is the delay time between the two-photon detections. In addition, $b_{BS}^{1}(t)=[b^{1}_{\rm out}(t)+b^{2}_{\rm out}(t)]/\sqrt{2}$, and $b_{BS}^{2}(t)=[b^{1}_{\rm out}(t)-b^{2}_{\rm out}(t)]/\sqrt{2}$, describe the output fields of photons after interfering at the BS [cf. Appendix C].
@@
In analogy to Eq. (25), the numerator of Eq. (27) corresponds to the area below the zeroth peak, and the normalization to the area of the high peak at the repetition time $\tau={\color[rgb]{0,0,0}1/R}$. The indistinguishability of two SPSs can be simply defined as $I=1-g^{(2)}_{\rm HOM}[0].$
We have computed $G_{\rm HOM}^{(2)}(\tau)$ as a function of $\tau$, for a setup with two emitters, parameters A of Table 1, and a train of 2 excitation pulses [cf. Fig. 5(b)]. Here, we observe a clear peak at the repetition time $\color[rgb]{0,0,0}\tau=1/R$} due to the detection of two nearly independent photons coming from two different pulses. The correlations $G_{\rm HOM}^{(2)}(\tau)$ are strongly suppressed at the origin, $\tau\sim 0$, showing only a minor zeroth peak which is enlarged in Fig. 5(c). The small area of this zeroth peak manifests the high indistinguishability of the single-photons. We also see that this behavior is nearly insensitive to the coupling $g/\kappa$, but may strongly depend on the decoherence parameters.
Figures 5(d)-(e) display $g^{(2)}_{\rm HOM}[0]$ as a function of the decoherence rates $\gamma_{\phi}/\gamma$ and $\gamma_{\rm loss}/\gamma.$ Note that in our implementation with parameters A we reach $I=0.986$, whereas for parameters B-D we obtain $I=0.993$, $I=0.998$, and $I=0.992$, respectively (cf. Table 1). We also find that the distinguishability $g^{(2)}_{\rm HOM}[0]$ is quite insensitive to $\gamma_{\rm loss}/\gamma$, but it dramatically increases with dephasing, following the analytical prediction $g^{(2)}_{\rm HOM}[0]\approx({\color[rgb]{0,0,0}2}\gamma_{\phi}/\gamma)/(1+{\color[rgb]{0,0,0}2}\gamma_{\phi}/\gamma)$ Bylander et al. (2003). This contrasts with the behaviour of $D_{2}$ in Fig. 3(e), illustrating that the dependence error $D_{2}$ of two SPSs is qualitatively different from the distinguishability $g^{(2)}_{\rm HOM}[0]$ measured by HOM interference. Consequently, both figures of merit have to be considered when designing high-performance multiplexed and synchronized SPSs.